# FinPos: A Position-Aware Trading Agent System for Real Financial Markets

### v1

@@ 6. Conclusion
In this paper, we introduce a position-aware trading task that more closely resembles real market conditions compared to the single-step trading task previously employed by researchers. To address this task, we design FinPos, a novel LLM Trading Agent equipped with position-awareness and risk management capabilities. To meet the demands of position management, an LLM agent must possess heightened market sensitivity, enhanced risk control abilities, and a longer-term perspective. Consequently, FinPos employs a specialized market perception module, a dual-decision agent system, and a multi-timescale reward design to achieve these objectives. Our experiments demonstrate that the position management task imposes higher comprehensive ability requirements on LLM Agents. FinPos explores solutions for trading agents operating in this more realistic market environment. However, we also identify two limitations of FinPos that require further attention. Firstly, FinPos is unable to implement coordinated management of multiple asset allocations, thus lacking the capability for higher-level operations such as risk hedging. Secondly, FinPos has been tested only on stock-type investment assets, and its applicability to other asset classes such as futures and bonds remains to be validated.
@@ Appendix A The implementation details of FinPos
A.1. Market Signal Processing and Analysis Module
@@
A.1.1. 10-K Filings (Annual Reports)
@@
A.1.2. 10-Q Filings (Quarterly Reports)
@@ 4. Evaluate the likely impact in both:
A.2. Dual Trading Decision Module
@@
A.2.1. Direction Decision Agent
@@ 2. Reflection:
A.2.2. Quantity and Risk Decision Agent
@@ Appendix B Formulas of Classic Financial Metrics
To better evaluate the risk–return profile of trading strategies, we adopt two widely used risk metrics: the Conditional Value at Risk (CVaR) and the Calmar Ratio. Their formal definitions are as follows.
B.1. Conditional Value at Risk (CVaR)
@@
In our experiments, $PnL$ is calculated based on realized returns over the testing period. A smaller CVaR indicates stronger downside protection and better risk control capability.
B.2. Calmar Ratio
@@
Appendix C More Experiments
@@
Figure 6. Profitability versus risk dynamics. Top-left: Calmar ratio, capturing return relative to maximum drawdown (higher is better). Bottom-left: joint view of profitability (CR%) and risk-control (1 - MDD), indicating return-risk balance. Right: time-series comparison of cumulative return (bottom) and exposure risk (top) across major events; PA-aware strategies mitigate exposure spikes and sustain more stable growth.
C.1. More Stock Trading Result Graphs
@@
C.2. More Details of Ablation Studies
C.2.1. Financial Insight Prompting (FIP)
This section explains the specific contents of the 8 dimensions in the fip section.(Tab. 3)
Table 3. Evaluation Criteria for Prompt Traits
@@ C.2.1. Financial Insight Prompting (FIP)
Role Setting Does it clearly assign a professional role to the AI (e.g., ”Financial Analyst”)?
@@
C.2.2. Macroeconomic News (MN) module

### v2

@@ 6 Conclusion
In this paper, we introduce a position-aware trading task that more closely resembles real market conditions compared to the single-step trading task previously employed by researchers. To address this task, we design FinPos, a novel LLM Trading Agent equipped with position-awareness and risk management capabilities. To meet the demands of position management, an LLM agent must possess heightened market sensitivity, enhanced risk control abilities, and a longer-term perspective. Consequently, FinPos employs a specialized market perception module, a dual-decision agent system, and a multi-timescale reward design to achieve these objectives. Our experiments demonstrate that the position management task imposes higher comprehensive ability requirements on LLM Agents. FinPos explores solutions for trading agents operating in this more realistic market environment.
Limitations
This work is intended for research purposes only. Deploying LLM-based trading systems in real-world financial markets without professional oversight may lead to financial losses, particularly under extreme or unexpected market conditions.
Single-Asset Focus
Although the proposed framework is general, in this paper we adopt a single-asset trading task as the experimental vehicle to validate its effectiveness.Extending the framework to portfolio-based methods and multi-asset allocation settings is an important direction for future work. This will involve addressing cross-asset correlation and rebalancing, their diversified risks, and portfolio-level dynamics, which presents additional complexities that require further model refinement.
Dependency on Prompt Quality:
major limitation of FinPOS lies in its sensitivity to prompt design. In our experiments, we focus on well-established companies with abundant and reliable textual information to ensure stable decision-making. As a result, the system relies on asset-specific prompt configurations that are tailored to the characteristics of individual stocks. This design choice limits the direct applicability of FinPOS to newly listed firms or assets with sparse or highly noisy textual signals, where prompt tuning and data curation become necessary. More broadly, this highlights a structural challenge of LLM-based trading agents: their performance is closely tied to the availability and organization of domain-specific textual inputs, rather than solely to model capacity. Addressing this limitation may ultimately require moving beyond prompt-level engineering. However, model-level adaptation or domain specialization typically demands substantial domain-specific data and computational resources, which we leave to future work.
Reward Design and Multi-Objective Optimization:
In this work, we adopt a lightweight multi-timescale reward design to regulate position risk and trend alignment, prioritizing stability, interpretability, and reproducibility. While reinforcement learning (RL) has shown promise for LLM-based agents, its application in financial markets is challenged by non-stationarity, heavy-tailed risks, and regime shifts, which can lead to overfitting and unstable behavior. Exploring principled RL formulations for learning general financial decision principles remains an important direction for future work.
@@ Appendix A The implementation details of FinPos
A.1 Market Signal Processing and Analysis Module
@@
A.1.1 10-K Filings (Annual Reports)
@@
A.1.2 10-Q Filings (Quarterly Reports)
@@ 4. Evaluate the likely impact in both:
A.2 Dual Trading Decision Module
@@
A.2.1 Direction Decision Agent
@@ 2. Reflection:
A.2.2 Quantity and Risk Decision Agent
@@ Appendix B Formulas of Classic Financial Metrics
To evaluate the risk-return characteristics of trading strategies, we summarize the formal definitions of commonly used financial evaluation metrics, including risk-adjusted return and downside risk measures, which are used throughout our experiments.
B.1 Definitions of Evaluation Metrics
$\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}},\quad$ (6)
$\text{MDD}=\max_{t=1}^{N}\left(\frac{P^{t}-P_{trough}^{t}}{P^{t}}\right)$ (7)
B.2 Conditional Value at Risk (CVaR)
@@
Importantly, during trading, CVaR is computed online using a 20-trading-day rolling window of past realized returns and updated daily, ensuring that no forward-looking or test-period information is used in position sizing. A smaller CVaR indicates stronger downside protection and more effective risk control.
B.3 Calmar Ratio
@@
Appendix C Experimental Setup
C.1 Position Awareness and Decision Structures
As summarized in Tab. 3, most existing LLM-based trading agents formulate decision-making as a discrete action selection problem, typically restricted to buy, sell, or hold, with fixed or implicit trade sizes. FinGPT adopts a single-agent architecture with predetermined position sizes, while FinMem leverages memory mechanisms to guide reasoning without explicitly modeling position magnitude or exposure. Although FinCon and FinAgent employ multi-stage or debate-based reasoning pipelines, their final outputs remain direction-only decisions without explicit position sizing. However, trading volume is a fundamental component of position awareness. Without the ability to adjust exposure magnitude, agents operating under direction-only actions lack fine-grained control over risk. For instance, when holding a large long position, an agent may detect increased downside risk but can only respond through unit-based actions, resulting in delayed or insufficient risk mitigation. In contrast, FinPos explicitly models position evolution through a two-stage decision structure that decouples directional reasoning from quantitative position sizing. By incorporating CVaR-based position control and reflection guided by multi-timescale rewards, FinPos enables risk-aware adjustments to both trading direction and exposure magnitude, leading to more coherent and realistic position management.
C.2 LLM-based Baselines Trading Agents
We compare FinPos against a representative set of state-of-the-art LLM-based trading agents that differ in architectural design, information processing pipelines, and decision structures, as summarized in Tab 3. We strictly follow the inference settings reported in their original papers.
Table 3: Overview of LLM-based Trading Agents and Their Architectures
Method Backbone Agents / Modules Position-Aware Information Processed Decision Structure
FinPOS GPT-4o 10 agents: 4 filters + 4 analyzers + Direction Agent + Quantity/Risk Agent Yes (continuous position state + CVaR sizing) Price, news, macro events, 10-K/10-Q reports Two-stage: (1) direction reasoning; (2) CVaR-based position sizing; reflection with multi-timescale reward
FinCon GPT-4o 7 agents: Data, News, Report, Analyst Report, Earnings Call, Stock Selection, Manager No (no explicit position sizing; only discrete actions) Structured & unstructured financial data Multistage pipeline with manager aggregation; buy/sell/hold
FinAgent GPT-4o 10+ agents: Fundamental Analyst, Sentiment Analyst, News Analyst, Technical Analyst, Bullish/Bearish Researchers, Trader, Risk (3 agents), Fund Manager No (no explicit position sizing; only discrete actions) Sentiment, technical indicators, news, fundamentals (e.g., revenue, profit, debt) Debate + risk-check pipeline; buy/sell/hold
FinGPT llama (fine-tuned) Single agent No Sentiment signals + price Generates buy/sell signals with fixed trade size
FinMem GPT-4o Single agent No (memory guides reasoning only) Data, News, 10-K/10-Q Single-step; buy/sell/hold
C.3 Deep Reinforcement Learning Baselines
In addition to LLM-based agents, we benchmark FinPos against classical deep reinforcement learning (DRL) agents, including A2C, PPO, and DQN. All DRL baselines are implemented using the FinRL framework Liu et al. (2020). These agents operate solely on numerical features derived from market prices and technical indicators, without access to textual or semantic information. The key hyperparameters for each RL agent are listed in Table 4.
Table 4: Key Hyperparameters for RL Agents
Agent Key Hyperparameters Value
A2C n_steps 5
ent_coef 0.01
learning_rate 0.0007
PPO n_steps 2048
ent_coef 0.01
learning_rate 0.00025
batch_size 64
DQN batch_size 128
buffer_size 50000
learning_rate 0.001
Figure 5: Prompt ablation across eight characteristics and three emphasis levels.
Appendix D More Experiments
@@
Figure 6: Profitability versus risk dynamics. Top-left: Calmar ratio, capturing return relative to maximum drawdown (higher is better). Bottom-left: joint view of profitability (CR%) and risk-control (1 - MDD), indicating return-risk balance. Right: time-series comparison of cumulative return (bottom) and exposure risk (top) across major events; PA-aware strategies mitigate exposure spikes and sustain more stable growth.
D.1 Sensitivity to LLM Sampling Hyperparameters
FinPOS does not train or fine-tune large language models; therefore, its performance is independent of random initialization seeds typically used in neural network training. However, to address reviewers’ concerns regarding stochasticity introduced during LLM inference, we conduct a sensitivity analysis by varying key decoding hyperparameters, including temperature and top_p.
Specifically, we evaluate FinPOS under multiple sampling configurations on TSLA while keeping all other components unchanged. As shown in Tab. 5, the overall performance remains stable across different settings, with only minor variations in cumulative return (CR), Sharpe ratio (SR), and maximum drawdown (MDD). These results indicate that FinPOS is not overly sensitive to reasonable changes in LLM sampling strategies, and its trading behavior is robust under inference-time stochasticity.
Table 5: Sensitivity Analysis of FinPOS under Different LLM Sampling Hyperparameters (TSLA)
Sampling Setting CR (%) SR MDD (%)
Default (TT=0.7, pp=0.9) 62.15 0.68 42.34
TT=0.7, pp=0.85 61.74 0.67 42.50
TT=0.7, pp=0.80 62.01 0.66 42.40
TT=0.8, pp=0.9 60.13 0.66 45.05
TT=0.5, pp=0.9 61.07 0.64 43.10
D.2 More Stock Trading Result Graphs
@@
D.3 More Details of Ablation Studies
D.3.1 Financial Insight Prompting (FIP)
Introducing FIP leads to a clear overall performance gain, consistently enhancing cumulative return across all assets (e.g., TSLA: 52.56% → 62.15%; AAPL: 59.38% → 67.31%). To investigate the impact of prompt design on the depth of agent finance insight, we divide the content of prompts into eight key dimensions: Role Setting, Task Breakdown, Depth Analysis, Multi-Dimensional Analysis, Logical Reasoning, Contextual Consideration, Actionability, and Clarity & Conciseness(the detailed explanations of these eight dimensions are provided in Tab. 6). For each dimension, we design prompts with different levels of refinement—Low, Medium, and High—and conduct graded experiments. The results are shown in Fig. 5.
Table 6: Evaluation Criteria for Prompt Traits
@@ D.3.1 Financial Insight Prompting (FIP)
Role Setting Does it clearly assign a professional role to the AI (e.g., "Financial Analyst")?
@@
D.3.2 Signal Ablation across Analyst Agents
To better isolate the contributions of individual information sources, we conduct an agent-wise signal ablation study by selectively disabling specific analyst agents while keeping all other components unchanged. This design allows us to quantify the marginal effect of each signal type under identical market conditions.
As this experiment aims to analyze the relative importance of different information sources rather than cross-asset generalization, we report representative results on TSLA, which offers the richest and most comprehensive set of textual signals. We observe consistent qualitative trends across other assets.
As shown in Table 7, removing any individual signal source leads to a noticeable degradation in performance, though the magnitude varies by signal type. Disabling quarterly filings (10-Q) results in a relatively moderate decline. In contrast, removing annual filings (10-K) substantially increases maximum drawdown, indicating their importance in anchoring long-term risk perception. Overall, company-level news has the most pronounced impact on both return and risk metrics, followed by macroeconomic signals.
Table 7: Agent-wise Signal Ablation on TSLA
Enabled Signals Performance
Market Company 10-Q 10-K CR% SR MDD%
62.15 0.68 42.34
$\times$ 51.32 0.53 55.40
$\times$ 48.15 0.48 60.52
$\times$ 58.59 0.63 45.11
$\times$ 53.78 0.56 67.25
D.3.3 Macroeconomic News (MN) module
@@
D.4 FinPOS Performance under Extreme Market Conditions
To evaluate the robustness of FinPOS under highly volatile market regimes, we conduct an additional stress test over the period from March to April 2025, during which market conditions exhibit pronounced instability. This interval coincides with a series of major macroeconomic and political events, including the U.S. presidential election and heightened geopolitical tensions, leading to frequent regime shifts and elevated tail risks.
Table 8 reports the performance of different methods during this period. DRL-based approaches suffer substantial drawdowns and consistently negative risk-adjusted returns, indicating limited robustness to abrupt distributional shifts. Similarly, most LLM-based agents operating under discrete buy/sell/hold decision frameworks display unstable behavior, particularly during rapid price reversals and high-volatility episodes. In contrast, FinPOS maintains strong and consistent performance across all evaluated stocks, achieving higher cumulative returns and Sharpe ratios while substantially reducing maximum drawdowns. Notably, the performance gap between FinPOS and baseline methods widens under extreme market conditions, where unmanaged position exposure can quickly amplify downside risk. We attribute this robustness to FinPOS’s explicit modeling of continuous position evolution and risk-aware sizing. By combining CVaR-based position control with multi-timescale reward feedback, FinPOS is able to gradually reduce exposure even when directional signals remain uncertain. This capability is especially critical during black-swan-like events.
Table 8: Extended results on TSLA, AAPL, and COIN (Mar–Sep 2025).
Models TSLA AAPL COIN
[…]
Baseline
Random 32.13-32.13 0.91-0.91 $62.10$ 34.21-34.21 0.28-0.28 $59.40$ 0.22-0.22 0.01-0.01 $17.60$
DRL
PPO Li et al. (2025) 73.76-73.76 0.56-0.56 $69.88$ 58.33-58.33 0.55-0.55 $61.40$ 2.04-2.04 0.04-0.04 $17.80$
LLM-Based Agents
FinGPT Yang et al. (2023) 95.19-95.19 1.07-1.07 $94.36$ 85.22-85.22 1.52-1.52 $74.60$ 3.63-3.63 0.04-0.04 $17.00$
FinAgent Zhang et al. (2024) 74.31-74.31 0.89-0.89 $85.65$ 78.00-78.00 1.09-1.09 $71.50$ 2.10-2.10 0.03-0.03 $17.50$
FINMEM Yu et al. (2024a) 44.03-44.03 0.52-0.52 $72.10$ 45.88-45.88 0.54-0.54 $47.80$ $0.13$ $0.01$ $19.20$
FINCON Yu et al. (2024b) $7.76$ $0.38$ $59.13$ 16.02-16.02 0.13-0.13 $29.03$ $7.35$ $1.03$ $15.10$
FinPos $54.99$ $0.67$ $42.34$ $60.28$ $0.69$ $19.75$ $14.74$ $0.92$ $14.05$